# ZALM3: Zero-Shot Enhancement of Vision-Language Alignment via In-Context Information in Multi-Turn Multimodal Medical Dialogue

### v1

@@ 5.1 Implementation Details
Although some Chinese multi-turn textual medical dialogue datasets are available [83, 84, 51, 85], to the best of our knowledge, there are no publicly accessible Chinese multi-turn multimodal medical dialogue datasets. Therefore, we comprehensively conduct experiments on our in-house datasets. Our data for the experiments includes the departments of dermatology, ophthalmology, and TCM, because these departments possess substantial amounts of image data as mentioned in Sec. 3.2. Moreover, the rate of poor image quality is high as shown in Fig. 2 (orange bars). For example, tongue diagnosis is an indispensable routine examination in TCM [86]. It plays a crucial role in the diagnosis and treatment of various diseases by combining tongue examination with clinical assessments. While a doctor requires a close-up of a patient’s tongue, the patient always takes a photo that includes her/his face, making the tongue region much smaller than what the doctor expects. A visualized example in Sec. 5.2.2 highlights that such vision-language misalignment downgrades the model’s response.
@@ 6 Discussion, Limitations, and Future Work
Many medical LLMs conducted performance comparisons with ChatGPT/GPT-4 or used them for scoring [24, 14, 15, 11]. Unfortunately, due to our platform’s medical data privacy protection policies, we are prohibited from calling external API interfaces including ChatGPT/GPT-4. However, those studies have already demonstrated that open-source LLMs trained with specialized medical data outperform general-purpose LLMs in professional fields. To achieve a better performance of keyword extractions in ZALM3, we believe a model fine-tuned with medical data can be adopted, such as the LLM part in our department-wise VLM. It is worth noting that the images sent by users can be regarded as “natural images with medical information” rather than typical medical images (DICOM grayscale format). While there are some outstanding medical multimodal representation models [89, 90, 91], a zero-shot visual grounding model trained on natural images, like GDINO, fits our scenario more.
@@
We observe that some patients upload photos of medications and lab reports, which contain extensive textual information, requiring the model to have the capacity of optical character recognition (OCR). With the recent emergence of studies on medical document question answering (DQA) [95], addressing this issue become important in our future work.

### v2

@@ 5.1 Implementation Details
Although some Chinese multi-turn textual medical dialogue datasets are available [83, 84, 51, 85], to the best of our knowledge, there are no publicly accessible Chinese multi-turn multimodal medical dialogue datasets. Therefore, we comprehensively conduct experiments on our in-house datasets. This study is reviewed and approved by the Ethics Committee of Qingdao Ping An Kangjian Internet Hospital (IRB number: LLSC2024A01).
Our data for the experiments includes the departments of dermatology, ophthalmology, and TCM, because these departments possess substantial amounts of image data as mentioned in Sec. 3.2. Moreover, the rate of poor image quality is high as shown in Fig. 2 (orange bars). For example, tongue diagnosis is an indispensable routine examination in TCM [86]. It plays a crucial role in the diagnosis and treatment of various diseases by combining tongue examination with clinical assessments. While a doctor requires a close-up of a patient’s tongue, the patient always takes a photo that includes her/his face, making the tongue region much smaller than what the doctor expects. A visualized example in Sec. 5.2.2 highlights that such vision-language misalignment downgrades the model’s response.
@@ 6 Discussion, Limitations, and Future Work
6.1 Clinical Relevance
Since the unforeseen COVID-19 outbreak, online doctor consultation applications have gradually become an integral component of clinical practice and continue to play an essential role in the post-pandemic era [14, chen2020exploring]. In developing countries, where healthcare resources are severely imbalanced, online consultations effectively enable patients in underdeveloped areas to access high-quality medical expertise. As the volume of online consultations grows, introducing AI-assisted manners becomes increasingly necessary to alleviate the overhead for human doctors. Recent research has primarily focused on text-only patient-doctor conversation scenarios. Unlike those studies, our scenario encompasses multimodal multi-turn medical consultations, incorporating both text and image modalities, which is a notable challenge in online doctor consultation applications that has yet to be well addressed. If responses to images cannot be generated effectively, it could severely harm the patient’s medical experience. For example in Fig. 3, the red content generated by the “original model” would significantly downgrade the patient’s satisfaction.
Our proposed ZALM3 has demonstrated its capacity to accurately identify and extract information from user-uploaded images, enhancing the quality of vision-language alignment during VLM training and inference. By enabling AI doctors to better understand patient descriptions of affected areas (through both text and images) and to more effectively interact with patients, ZALM3 can significantly improve diagnostic efficiency and accuracy in real clinical settings, further reducing the labor costs for human doctors.
6.2 Privacy Constraints
Many medical LLMs conducted performance comparisons with ChatGPT/GPT-4 or used them for scoring [24, 14, 15, 11]. Unfortunately, due to our platform’s medical data privacy protection policies, we are prohibited from calling external API interfaces including ChatGPT/GPT-4. However, those studies have already demonstrated that open-source LLMs trained with specialized medical data outperform general-purpose LLMs in professional fields. Given the stringent data privacy requirements in medical environments, ZALM3’s compliance with data protection policies ensures it can safely be integrated into clinical practice without compromising sensitive patient information.
6.3 Model Selection in ZALM3
To achieve a better performance of keyword extractions in ZALM3, we believe a model fine-tuned with medical data can be adopted, such as the LLM part in our department-wise VLM. It is worth noting that the images sent by users can be regarded as “natural images with medical information” rather than typical medical images (DICOM grayscale format). While there are some outstanding medical multimodal representation models [89, 90, 91], a zero-shot visual grounding model trained on natural images, like GDINO, fits our scenario more.
6.4 Limitations
@@
6.5 Future Work
We observe that some patients upload photos of medications and lab reports, which contain extensive textual information, requiring the model to have the capacity of optical character recognition (OCR). With the recent emergence of studies on medical document question answering (DQA) [95], addressing this issue becomes important in our future work.